OpenStudy (anonymous):

The area of a rectangle is 51 square inches. If the width of the rectangle is 6 inches, what is the length? 19.5 inches 13 inches 9 inches 8.5 inches

OpenStudy (hoblos):

area of rectangle = l*w if area = 51 square inches and width is 6 inches l= area/width = 51/6 = 8.5inches